# GeoRecon: Graph-Level Representation Learning for 3D Molecules via Reconstruction-Based Pretraining

### v1

@@ Abstract
The pretraining-and-finetuning paradigm has driven significant advances across domains, such as natural language processing and computer vision, with representative pretraining paradigms such as masked language modeling and next-token prediction. However, in molecular representation learning, the task design remains largely limited to node-level denoising, which is effective at modeling local atomic environments, yet maybe insufficient for capturing the global molecular structure required by graph-level property prediction tasks, such as energy estimation and molecular regression. In this work, we present GeoRecon, a novel graph-level pretraining framework that shifts the focus from individual atoms to the molecule as an integrated whole. GeoRecon introduces a graph-level reconstruction task: during pretraining, the model is trained to generate an informative graph representation capable of accurately guiding reconstruction of the molecular geometry. This encourages the model to learn coherent, global structural features rather than isolated atomic details. Without relying on additional supervision or external data, GeoRecon outperforms node-centric baselines on multiple molecular benchmarks (e.g., QM9, MD17), demonstrating the benefit of incorporating graph-level reconstruction for learning more holistic and geometry-aware molecular embeddings.
@@ 1 Introduction
Predicting molecular properties from three-dimensional structures is central to modern chemistry and materials science. While Density Functional Theory (Parr, 1983) delivers high-fidelity estimates, its steep computational cost prevents large-scale virtual screening. This bottleneck has driven the development of numerous deep learning approaches that aim to predict molecular properties with comparable accuracy while offering significant speed advantages (Thölke and De Fabritiis, 2022; Li et al., 2023, 2024; Gasteiger et al., 2020a, 2021; Batatia et al., 2022). Concurrently, several studies (Zaidi et al., 2023; Feng et al., 2023; Fang et al., 2022; Zhou et al., 2023; Ji et al., 2024) have adapted the pretraining-and-finetuning paradigm—originally established in computer vision (Girshick et al., 2014) and natural language processing (Radford et al., 2018)—to deep learning for molecular modeling, demonstrating effective transfer of knowledge from unlabeled molecular datasets to downstream tasks.
To date, most molecular pre-training focuses on node denoising (Zaidi et al., 2023; Feng et al., 2023) or masked atom prediction (Hou et al., 2022; Xia et al., 2023), which effectively capture local atomic interactions but lack explicit mechanisms to encode holistic molecular geometry. In practice, many target properties, such as energy and dipole moment (Ramakrishnan et al., 2014; Chmiela et al., 2017), are determined by global structure rather than local interactions. This inherent difference could lead to representation incompatibility, potentially diminishing the benefits of pre-training for global-level downstream tasks (Hu et al., 2019). To overcome these limitations, recent methods incorporate explicit supervision via graph-level attributes (Hu et al., 2019) or frequent functional motifs (Rong et al., 2020), or adopt contrastive learning (Liu et al., 2021a). Nevertheless, these approaches are often either over specialized by hand-crafted supervisions to be broadly applicable, or are primarily designed for 2D or 2D&3D graphs, lacking inherent support for 3D molecular pre-training like denoising (Zaidi et al., 2023).
In this work, we introduce a graph-level 3D self-supervised pre-training task, which we call Geometric Reconstruction. Specifically, given a 3D molecule, the model is asked to generate an informative, invariant graph-level representation, which can be used to reconstruct the molecule’s entire 3D geometry. Since the graph-level representation is invariant to orientation, relying solely on it for reconstruction can lead to arbitrary orientations, making supervision challenging. To address this, we frame the reconstruction as a more demanding denoising task (Zaidi et al., 2023)—applying larger noise but using a lighter denoiser (decoder)—while conditioning the denoising process on the graph-level representation for guidance. This paradigm requires the model to learn a highly informative graph-level representation to compensate for the large individual noise. See Figure 1 for an illustration. Compared to directly supervising on some graph-level properties, this design is generic and task-agnostic.
To demonstrate the effectiveness of this task, we fine-tune the pre-trained model on widely adopted benchmarks, QM9 (Ramakrishnan et al., 2014) and MD17 (Chmiela et al., 2017), both of which focus on graph-level molecular property prediction. Our evaluations show that this straightforward extension generally improves the standard denoising baseline (Zaidi et al., 2023) and matches the performance of more sophisticated models that rely on external data, such as molecular spectra (Wang et al., 2025). This demonstrates that explicitly modeling graph-level structure enhances both accuracy and robustness in molecular representation learning. Additionally, we conduct several ablation studies on key design choices—such as lightweight decoder depth and noise scale in the reconstruction task—further confirming the robustness of our design.
2 Preliminaries
@@
$\mathbb{E}_{q_{\sigma}(\widetilde{\bm{r}},\bm{r})}\left[\left\|\mathrm{GNN}_{\theta}(\widetilde{\bm{r}})-\nabla_{\widetilde{\bm{r}}}\log q_{\sigma}(\widetilde{\bm{r}}\mid\bm{r})\right\|^{2}\right]=\mathbb{E}_{q_{\sigma}(\widetilde{\bm{r}},\bm{r})}\left[\left\|\mathrm{GNN}_{\theta}(\widetilde{\bm{r}})-\frac{\bm{r}-\widetilde{\bm{r}}}{\sigma^{2}}\right\|^{2}\right].$
$\bm{r}$ denotes the equilibrium molecular structure, $\widetilde{\bm{r}}$ is its perturbed version obtained by Gaussian corruption, and $\theta$ represents the parameters of the denoising network. $q_{\sigma}(\widetilde{\bm{r}}\mid\bm{r})$ is defined as a Gaussian centered at $\bm{r}$, serving as a tractable approximation to the Boltzmann distribution $p_{\text{physics}}(\widetilde{\bm{r}})\propto\exp(-E(\widetilde{\bm{r}}))$. Its score function corresponds to the standardized noise vector $(\bm{r}-\widetilde{\bm{r}})/\sigma^{2}$. More broadly, $p_{\text{physics}}$ can be approximated by a Gaussian mixture $q_{\sigma}(\widetilde{\bm{r}})\sim\frac{1}{n}\sum_{i}\mathcal{N}(\widetilde{\bm{r}};\bm{r}_{i},\sigma^{2}I)$, whose score defines a force field over noisy conformations.
@@
Building upon coordinate denoising, Feng et al. (2023) introduced additional dihedral angle noise $\delta\psi\sim\mathcal{N}(0,\sigma^{2}I_{m})$ along rotatable bonds. For small angular displacements, these perturbations can be approximately linearized and translated into structured coordinate noise $\bm{\epsilon}_{a}\sim\mathcal{N}(0,\sigma^{2}\bm{C}\bm{C}^{\top})$, where $\bm{C}$ denotes the transformation matrix from angular displacements to Cartesian coordinates. The resulting intermediate coordinates are given by $\bm{r}_{a}:=\bm{r}+\bm{\epsilon}_{a}$. A second isotropic coordinate perturbation $\bm{\epsilon}_{cd}\sim\mathcal{N}(0,\tau^{2}I_{3N})$ is then added, and the model is trained to predict this final perturbation $\bm{\epsilon}_{cd}$ during pretraining.
3 Method
@@
While Feng et al. (2023) advance coordinate denoising by incorporating molecular flexibility through dihedral perturbations, their framework still focuses on node-level supervision. The model predicts local coordinate noise for each atom, but without explicit mechanisms to capture global structural context, which limits its effectiveness on graph-level tasks.
@@
Building on the self-supervised, physically grounded coordinate-denoising objective (Zaidi et al., 2023; Feng et al., 2023), we exploit its dual benefit as both a force-field surrogate and a fine-grained geometric bias. Yet, when used in isolation, this purely node-level signal cannot enforce the global structural coherence required by graph-level property prediction.
@@
We therefore introduce GeoRecon, a reconstruction-based pretraining framework designed to capture global molecular structure from noisy inputs. Instead of directly decoding coordinates from pooled embeddings—which often lose spatial granularity—we formulate reconstruction as a denoising process over entire molecular geometries. Specifically, we inject stronger gaussian noise into the 3D coordinates and predicting the added noise with a lightweight decoder conditioned on graph-level representations.
@@
More precisely, GeoRecon builds upon standard coordinate denoising, but enhances its capacity to model global molecular structure by increasing the difficulty of the denoising task and providing graph-level representations as additional guidance. By injecting stronger, orientation-invariant noise into the coordinates and conditioning the decoder on a global embedding derived from clean structures, we encourage the encoder to attend to long-range dependencies and capture coherent, molecule-level geometry. This training scheme allows GeoRecon to extract more transferable representations, bridging local interactions with global structural awareness.
Notations. To support multi-task supervision, we construct three coordinate variants per molecule:
@@
  • pos_clean: the original coordinates, serving as the ground truth for the denoising task and simultaneously as the input for the reconstruction encoder task introduced in the following section, denoted by $\bm{r}_{i}^{\texttt{clean}}\in\mathbb{R}^{3}$, where $i$ indexes the atom in the molecule;
  • pos_noised: the perturbed atomic coordinates used as input for the standard denoising task, denoted by $\bm{r}_{i}^{\texttt{noised}}\in\mathbb{R}^{3}$;
  • pos_rec: a noise-scaled variant of the perturbed coordinates, used as the target in the reconstruction task; denoted by $\bm{r}_{i}^{\texttt{rec}}\in\mathbb{R}^{3}$.
@@
where $\lambda\in\mathbb{(}1,+\infty)$ modulates the reconstruction difficulty. Each target coordinate is paired with its corresponding atomic number $z_{i}$ to construct the initial atomic embeddings.
[…]
Pairwise edge features and edge embeddings are denoted as $\bm{f}_{ij}$ and $\mathbf{u}_{ij}^{(*)}$ respectively. For simplicity, we construct the rough edge feature $\mathbf{d}_{ij}$ between two atoms $i$ and $j$ by their relative positions and define: $\bm{\mathrm{d}}_{ij}=\bm{r}_{j}^{*}-\bm{r}_{i}^{*},\quad\hat{\bm{d}}_{ij}={\bm{\mathrm{d}}_{ij}}/{\|\bm{\mathrm{d}}_{ij}\|}$.
The model is trained to predict noise vectors $\hat{\bm{\epsilon}}_{i}^{(*)}\in\mathbb{R}^{3}$ for each atom and task, where the supervision signal is the synthetic Gaussian noise $\bm{\epsilon}_{i}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$ applied during pretraining.
We use $\mathcal{L}_{\text{denoise}}$, $\mathcal{L}_{\text{rec}}$, and $\mathcal{L}_{\text{clean}}$ to denote the three loss terms, with weighting coefficients $\lambda_{\text{denoise}}$, $\lambda_{\text{rec}}$, and $\lambda_{\text{clean}}$, respectively.

### v2

@@ Abstract
The pretraining–finetuning paradigm has powered major advances in domains such as natural language processing and computer vision, with representative examples including masked language modeling and next-token prediction. In molecular representation learning, however, pretraining tasks remain largely restricted to node-level denoising, which effectively captures local atomic environments but is often insufficient for encoding the global molecular structure critical to graph-level property prediction tasks such as energy estimation and molecular regression. To address this gap, we introduce GeoRecon, a graph-level pretraining framework that shifts the focus from individual atoms to the molecule as an integrated whole. GeoRecon formulates a graph-level reconstruction task: during pretraining, the model is trained to produce an informative graph representation that guides geometry reconstruction while inducing smoother and more transferable latent spaces. This encourages the learning of coherent, global structural features beyond isolated atomic details. Without relying on external supervision, GeoRecon achieves generally improves over backbones baselines on multiple molecular benchmarks including QM9, MD17, MD22, and 3BPA, demonstrating the effectiveness of graph-level reconstruction for holistic and geometry-aware molecular embeddings.
GeoRecon: Graph-Level Representation Learning for 3D Molecules via Reconstruction-Based Pretraining
Shaoheng Yan1,2          Zian Li1,3          Muhan Zhang1,∗
1Institute for Artificial Intelligence, Peking University 2Yuanpei College, Peking University 3School of Intelligence Science and Technology, Peking University
@@ 1 Introduction
With the pretraining–finetuning paradigm extending from computer vision and natural language processing  (Bao et al., 2021; Lu et al., 2019) to an increasingly broad range of domains, designing effective pretraining tasks that enable models to achieve competitive downstream performance with less data and faster convergence has become a key component in training pipeline design. In molecular representation learning, pretraining strategies based on denoising or masking have been widely validated for their effectiveness (Zaidi et al., 2023; Zhou et al., 2023; Feng et al., 2023). In these tasks, molecular structures are perturbed or masked at the atomic (node) level, and the model is trained to recover the original data.
However, a substantial portion of downstream tasks in molecular representation learning require the model to possess a strong understanding at the whole-molecule (graph) level (Li et al., 2020). Tasks such as energy or interatomic force prediction demand not only that the learned representation manifold faithfully capture node-level properties, but also that the model encode emergent graph-level information into the molecular representation. Unfortunately, conventional node-level pretraining tasks, by their very design, often lack explicit alignment with graph-level objectives, thereby limiting the potential gains that can be realized during the pretraining stage, motivating a shift from purely node-level perturbation to graph-level supervision.
We conducted a preliminary experiment by perturbing the atomic coordinates of a molecule with small displacements and using a spectral norm method described in Appendix B to estimate the Lipschitz constant of the encoder in the public Coord (Zaidi et al., 2023) pretraining checkpoint, a standard node-denoising-based baseline. This constant measures the sensitivity of the encoder’s output representation to coordinate changes (larger values indicate higher sensitivity).
Table 1: Local non-rigid Lipschitz constant $L(x)$ for GeoRecon and the Coord baseline under two parameter settings. steps is the number of iterations in the power method used to estimate the dominant singular value.
Model step 5 15 25
GeoRecon median $2.969\text{\times}{10}^{1}$ $3.071\text{\times}{10}^{1}$ $3.075\text{\times}{10}^{1}$
p95 $5.087\text{\times}{10}^{1}$ $5.090\text{\times}{10}^{1}$ $5.090\text{\times}{10}^{1}$
Coord median $2.537\text{\times}{10}^{4}$ $2.539\text{\times}{10}^{4}$ $2.539\text{\times}{10}^{4}$
p95 $4.629\text{\times}{10}^{4}$ $638\text{\times}{10}^{4} 4.4$$451\text{\times}{10}^{4}$
As shown in the second row of Table 1, Coord exhibits values exceeding $10^{4}$, indicating extreme sensitivity to minute perturbations. This instability underscores the limitations of node-level pretraining, as smooth representation manifolds are well known to facilitate downstream performance (Lee et al., 2025; Guo et al., 2024; Zhang et al., 2025; Krishnan et al., 2020). We provide a brief proof in Appendix C.
To address such limitations, prior work has introduced explicit supervision via graph-level attributes (Hu et al., 2019) or frequent functional motifs (Rong et al., 2020), or adopted contrastive learning (Liu et al., 2021a). However, these methods are often over-specialized by hand-crafted supervision, limiting their general applicability, or are primarily designed for 2D or mixed 2D&3D settings, lacking the inherent support for fully 3D molecular pretraining offered by denoising-based frameworks (Zaidi et al., 2023).
To further bridge the gap between node and graph-level pretraining, we propose Geometric Reconstruction, a graph-level 3D self-supervised pretraining task. Given a 3D molecular structure, the model is trained to produce an informative, orientation-invariant graph-level representation that can guide the reconstruction of the molecule’s full 3D geometry from heavily noised coordinates. This forces the encoder to capture rich, molecule-level structural information sufficient to recover heavily perturbed geometries.
GeoRecon possesses several key properties that make it distinct from prior graph-level pretraining approaches and broadly applicable across molecular modeling scenarios:
  • 3D-only input: Relies solely on atomic coordinates, without requiring 2D bond graphs or structural annotations, avoiding dependence on noisy or ambiguous 2D labels;
  • Fully self-supervised: Learns directly from raw coordinates without any external labels;
  • Label-free and broadly applicable: Does not require additional annotations (e.g., spectral data), enabling application to a wide range of molecular datasets;
  • Architecture-agnostic pretraining objective: Can be integrated into different SE(3)-equivariant backbones without altering downstream fine-tuning protocols.
To demonstrate the effectiveness of GeoRecon, we fine-tune the pre-trained model on a range of graph-level molecular property prediction benchmarks, including QM9 (Ramakrishnan et al., 2014), MD17 (Chmiela et al., 2017), and the challenging MD22 (Chmiela et al., 2023) dataset. Our method consistently outperforms our backbone on MD17 and MD22, and achieves improvements on most metrics of QM9.
We further assess the robustness and generality of GeoRecon through two parallel evaluations: (i) ablations on training set size, encoder depth, pooling strategy, decoder depth, and reconstruction noise scale, and (ii) an out-of-distribution evaluation on 3BPA comparing GeoRecon with its TorchMD-Net backbone. Taken together, these results indicate that incorporating explicit graph-level structure in pretraining enhances both accuracy and robustness, with particularly strong benefits for challenging molecular modeling tasks that depend heavily on graph-level information.
2 Related Works
Implicit Use of 3D Information Some prior works implicitly leverage 3D structural information by converting 1D SMILES strings or 2D molecular graphs into 3D conformations using cheminformatics tools such as RDKit (Liu et al., 2021a; Zhu et al., 2022; Stärk et al., 2022). These derived conformers are then used during training to enhance model capacity, achieving strong results on benchmarks like PCQM4Mv2 (Hussain et al., 2024), which are originally designed for 2D graph modalities.
Self‑Supervised Denoising Pretraining Self‑supervised denoising has emerged as a principled means to leverage unlabeled molecular conformations. Coord pretraining (Zaidi et al., 2023) demonstrates that training a GNN to remove Gaussian perturbations from equilibrium structures is mathematically equivalent to learning the underlying molecular force field, markedly improving downstream QM9 performance under limited labels. Frad (Feng et al., 2023) furthers this concept by jointly perturbing dihedral angles and Cartesian coordinates, better emulating realistic conformational variations. SliDe (Ni et al., 2023) augments denoising tasks with classical force‑field priors (bond lengths, angles, torsions), grounding the learned representation in established physical interactions. More recently, MolSpectra (Wang et al., 2025) integrates multimodal energy‑level spectra via contrastive reconstruction, enriching the model’s access to discrete quantum information without external labels.
Multi-Task and Multi-Fidelity Learning To address data scarcity in quantum chemistry, one strategy is to exploit shared structure across tasks or fidelity levels. In multi-task learning, models predict several quantum properties simultaneously (e.g., 12 QM9 targets), encouraging a shared embedding space and reducing overfitting. In multi-fidelity learning, one leverages the hierarchy of computational costs: accurate methods like CCSD(T) are expensive, while cheaper ones like DFT scale to larger datasets. Recent work (Ren et al., 2023) combines abundant DFT data with limited CC samples to approach CC-level accuracy at far lower cost. Together, these strategies mitigate scarcity by balancing accuracy and efficiency.
Graph-Level Pretraining in Molecular Representation Learning Early GNN pretraining mainly targeted node-level tasks such as masked atom prediction (Hu et al., 2019; Hou et al., 2022; Xia et al., 2023), but recent efforts increasingly adopt graph-level objectives to capture molecule-wide signals. GROVER (Rong et al., 2020) introduces motif prediction to identify recurring substructures, and MGSSL (Zhang et al., 2021) extends this by generating graphs motif-by-motif, injecting global inductive biases. GraphCL (You et al., 2020) and GCC (Qiu et al., 2020) apply graph-level contrastive learning with augmented graph views to enforce whole-graph consistency.
Other approaches leverage cross-modal or multi-scale alignment: GraphMVP (Liu et al., 2021a) aligns 2D graphs with 3D conformers via contrastive learning, while Uni-Mol2 (Ji et al., 2024) uses a two-track transformer jointly encoding atomic, graph, and spatial features. Though effective, these methods often require extra modalities, augmentation pipelines, or complex designs. In contrast, our method employs a single graph-level reconstruction objective—conditioned on pooled global embeddings—built directly atop coordinate denoising, achieving strong downstream performance without auxiliary data or heavy overhead.
3 Preliminary
@@
$\mathbb{E}_{q_{\sigma}(\widetilde{\bm{r}},\bm{r})}$ $\left[\left\|\mathrm{GNN}_{\theta}(\widetilde{\bm{r}})-\nabla_{\widetilde{\bm{r}}}\log q_{\sigma}(\widetilde{\bm{r}}\mid\bm{r})\right\|^{2}\right]$
$=\mathbb{E}_{q_{\sigma}(\widetilde{\bm{r}},\bm{r})}\left[\left\|\mathrm{GNN}_{\theta}(\widetilde{\bm{r}})-\frac{\bm{r}-\widetilde{\bm{r}}}{\sigma^{2}}\right\|^{2}\right].$
$\bm{r}$ denotes the equilibrium molecular structure, $\widetilde{\bm{r}}$ is its perturbed version obtained by Gaussian corruption, and $\theta$ represents the parameters of the denoising network. $q_{\sigma}(\widetilde{\bm{r}}\mid\bm{r})$ is defined as a Gaussian centered at $\bm{r}$, serving as a tractable approximation to the Boltzmann distribution $)p_{\text{physics}}(\widetilde{\bm{r}})\propto\exp(-E(\widetilde{\bm{r}}$)). Its score function corresponds to the standardized noise vector $(\bm{r}-\widetilde{\bm{r}})/\sigma^{2}$. More broadly, $p_{\text{physics}}$ can be approximated by a Gaussian mixture $q_{\sigma}(\widetilde{\bm{r}})\sim\frac{1}{n}\sum_{i}\mathcal{N}(\widetilde{\bm{r}};\bm{r}_{i},\sigma^{2}I)$, whose score defines a force field over noisy conformations.
@@
Building upon coordinate denoising, Feng et al. (2023) introduced additional dihedral angle noise $\delta\psi\sim\mathcal{N}(0,\sigma^{2}I_{m})$ along rotatable bonds. While Feng et al. (2023) advance coordinate denoising by incorporating molecular flexibility through dihedral perturbations, their framework still focuses on node-level supervision. The model predicts local coordinate noise for each atom, but without explicit mechanisms to capture global structural context, which limits its effectiveness on graph-level tasks.
4 Our Method
Figure 1: Overview of the GeoRecon framework. Given a molecular structure with atom types and 3D coordinates, the model encodes it using SE(3)-equivariant attention. Besides the standard node-level denoising objective, GeoRecon feeds a pooled graph-level representation concatenated with node embeddings derived from noisy coordinates into a lightweight decoder to reconstruct the scaled noise. The pretrained encoder is then finetuned for downstream molecular property prediction tasks.
@@ 4 Our Method
Building on the self-supervised, physically grounded coordinate-denoising objective (Zaidi et al., 2023; Feng et al., 2023), we exploit its dual benefit as both a force-field surrogate and a fine-grained geometric bias. Yet, when used in isolation, this purely node-level signal cannot enforce the global structural coherence required by graph-level property prediction. This gap motivates the need for a complementary mechanism that explicitly encourages the encoder to capture molecule-wide dependencies rather than only local perturbations.
@@
We therefore introduce GeoRecon, a reconstruction-based pretraining framework designed to capture global molecular structure from noisy inputs. A naive way is to directly decode coordinates from pooled graph-level embeddings. However, this method often loses spatial granularity. Furthermore, predicting equivariant coordinates from invariant graph embeddings means overfitting to specific orientations. Therefore, we formulate reconstruction as a denoising process too over entire molecular geometries. Specifically, we inject scaled Gaussian noise into the 3D coordinates and train a lightweight decoder, conditioned on graph-level representations, to predict the scaled noise.
@@
More precisely, GeoRecon extends standard coordinate denoising by increasing the task difficulty and conditioning the decoder on graph-level representations. Unlike approaches that carefully adjust noise to match specific physical semantics (Feng et al., 2023; Liu et al., 2025), we deliberately add orientation-invariant noise into the coordinates and reduce the layer-num of the decoder. This places greater reliance on the global embedding, compelling the encoder to capture long-range dependencies and coherent molecule-level geometry. As a result, GeoRecon learns more transferable representations that bridge local interactions with global structural awareness.
4.1 Notations
@@
To support multi-task supervision, we construct three coordinate variants per molecule. (i) Pos_clean: the original coordinates, serving as the ground truth for the denoising task and simultaneously as the input for the reconstruction encoder task introduced in the following section, denoted by $\bm{r}_{i}^{\texttt{cln}}\in\mathbb{R}^{3}$, where $i$ indexes the atom in the molecule; (ii) Pos_noised: the perturbed atomic coordinates used as input for the standard node-level denoising task, denoted by $\bm{r}_{i}^{\texttt{nsd}}\in\mathbb{R}^{3}$; (iii) Pos_rec: coordinates perturbed with an alternative noise scale $\lambda$, serving as the reconstruction target; denoted by $\bm{r}_{i}^{\texttt{rec}}\in\mathbb{R}^{3}$.
@@
$\bm{r}_{i}^{\texttt{nsd}}=\bm{r}_{i}^{\texttt{cln}}+\bm{\epsilon},\quad\bm{r}_{i}^{\texttt{rec}}=\bm{r}_{i}^{\texttt{cln}}+\lambda\cdot\bm{\epsilon}$
where $\lambda$ modulates the reconstruction difficulty, set to 1 for default. Each target coordinate is paired with its corresponding atomic number $z_{i}$ to construct the initial atomic embeddings.
@@
The model is trained to predict noise vectors $\hat{\bm{\epsilon}}_{i}^{(*)}\in\mathbb{R}^{3}$ for each atom and task, where the supervision signal is the synthetic Gaussian noise applied during pretraining. We denote the loss terms by $\mathcal{L}_{*}$, where the subscript indicates the task name, and use $\lambda_{*}$ for the corresponding weighting coefficients.